# Legged Robot State Estimation Using Invariant Neural-Augmented Kalman Filter with a Neural Compensator

### v1

@@ Abstract
This paper presents an algorithm to improve state estimation for legged robots. Among existing model-based state estimation methods for legged robots, the contact-aided invariant extended Kalman filter defines the state on a Lie group to preserve invariance, thereby significantly accelerating convergence. It achieves more accurate state estimation by leveraging contact information as measurements for the update step. However, when the model exhibits strong nonlinearity, the estimation accuracy decreases. Such nonlinearities can cause initial errors to accumulate and lead to large drifts over time. To address this issue, we propose compensating for errors by augmenting the Kalman filter with an artificial neural network serving as a nonlinear function approximator. Furthermore, we design this neural network to respect the Lie group structure to ensure invariance, resulting in our proposed Invariant Neural-Augmented Kalman Filter (InNKF). The proposed algorithm offers improved state estimation performance by combining the strengths of model-based and learning-based approaches. Supplementary Video: https://youtu.be/k1ZVb6Xj8D8
@@ III-A SE2(3) Group Generation Network
$\textbf{X}_{t}=\begin{bmatrix}\textbf{R}_{t}&v_{t}&p_{t}\\ \textbf{0}_{3,1}&1&0\\ \textbf{0}_{3,1}&0&1\\ \end{bmatrix}\in\text{SE}_{2}(3)$ (9)
@@
Using the linear combination of the generators, the neural compensator, as shown in Figure 1, consists of three main steps. First, a temporal convolutional network (TCN) [18] takes as input a time-series sequence that includes linear acceleration, angular velocity, joint position, joint velocity, foot position, foot velocity, contact state, and the estimated state from the InEKF over the previous 49 time-steps as well as the current time step, and it produces 9 outputs. Second, these outputs are combined linearly with the generators of $\mathfrak{se}_{2}$(3) to obtain an element of $\mathfrak{se}_{2}$(3), denoted as $\hat{\eta}_{t}^{r}$. Finally, by applying the exponential mapping described in (11), the result is mapped to an element of the SE2(3) group, $\hat{\textbf{E}}_{t}^{r}$, which represents the estimated error. This value acts as a compensation term for the estimated state in the InEKF.
@@ III-B Invariant Neural-Augmented Kalman Filter
The Neural Compensator introduced in Section III-A is augmented into the InEKF [5] to form the Invariant Neural-Augmented Kalman Filter (InNKF). The structure of InNKF is shown in Figure 2, where the state estimation is first performed using a model-based approach. In this process, two types of states are defined. The state of the InEKF is expressed as $\textbf{Z}_{t}\in\text{SE}_{N+2}(3)$, as shown in (12), while the state of the InNKF is expressed as $\textbf{X}_{t}\in\text{SE}_{2}(3)$. Here, $\textbf{X}_{t}$ is a $5\times 5$ submatrix in the top-left corner of $\textbf{Z}_{t}$.
@@
Using the SE2(3) Group Generation Network (SEGGN) designed in Section III-A, a Neural Compensator (NC) is added. The NC is trained on a dataset obtained from a simulation in which a single trajectory traverses four different terrains (stairs, slope, random uniform, discrete obstacle) for 100 seconds. The dataset included the ground truth state $\textbf{X}_{t}$, sensor data such as linear acceleration ($\tilde{a}_{t}$), angular velocity ($\tilde{\omega}_{t}$), joint position and velocity ($\tilde{q}t,\tilde{\dot{q}}t$), foot position and velocity ($\tilde{\textbf{p}}_{WC_{i}}(t),\tilde{\textbf{v}}_{WC_{i}}(t)$), contact state ($\text{S}_{t}$), and the estimated state $\bar{\textbf{X}}_{t}^{+}$ from the InEKF. This dataset was used to train the SEGGN through the supervised learning process shown in Figure 3.
@@ IV-A Performance Analysis of SE2(3) Group Generation
The neural network used in this study does not perform operations directly on the Lie group but operates in the Lie algebra, followed by exponential mapping. Since the Lie group is defined on a smooth manifold, performing group operations requires matrix multiplication. In contrast, the Lie algebra, defined in the tangent space, only requires matrix addition for operations. Consequently, obtaining the output in the Lie algebra and applying exponential mapping significantly improves training efficiency. Table I and Figure 4 show the computation time according to batch size. While matrix addition has a time complexity of O$(n^{2})$, matrix multiplication has a time complexity of O$(n^{3})$. This highlights the advantage of obtaining outputs in the tangent space, where only matrix addition is required, to enhance training efficiency. A difference of approximately 0.1 times in computation time was observed, confirming significantly faster performance.
Furthermore, due to the structure of the SEGGN, which utilizes the bases of $\mathfrak{se}_{2}(3)$ followed by an exponential mapping process, it inevitably produces outputs within SE2(3). The corresponding results are presented in Figure 5. The first row of Figure 5 illustrates the transformation of the original coordinates based on the output values obtained from SEGGN. Visually, it can be observed that the transformed coordinates maintain orthonormality, ensuring the structural integrity of the transformation. A numerical analysis of these results is presented in the histograms below. The Frobenius norm error distribution shows that nearly all samples have errors close to zero, indicating minimal deviation from the expected transformation properties. Additionally, the determinant values are consistently close to one, confirming that the rotational component of the output matrix adheres well to the properties of the SO(3) group (26).
@@ IV-B Implementation Details
Figure 6: Position estimation results across four terrains (stairs, slope, uneven terrain, discrete obstacle) over a time interval of 0 $\sim$ 140 s. The solid black line represents the ground truth, the blue dashed line represents InEKF, and the red dashed line represents the InNKF (proposed method).
To reduce the sim-to-real gap, sensor noise and IMU bias were introduced to better reflect real-world conditions. These noise factors were applied during both locomotion controller training and dataset collection. The sensor noise was assumed to follow a Gaussian distribution. Additionally, to address uncertainties in the measurement model, a covariance matrix was configured. Since InEKF measurements primarily depend on kinematic information, the covariance matrix was constructed based on the standard deviation of the noise applied to joint position and velocity.
@@ IV-C Experimental Results
Two baselines were compared to analyze the experimental results. The first was the model-based (MB) estimator, represented by the Invariant Extended Kalman Filter (InEKF). The second was the learning-based (LB) estimator, represented by a Neural Network (NN).
The training and test datasets were collected in the Isaac Gym simulator [19], using terrains shown in Figure 7. The quadruped robot, Unitree Go1, was the model used, controlled by a blind locomotion controller trained via reinforcement learning [20]. To evaluate whether the state estimator overfits the training conditions, additional experiments were conducted in Raisim [21]. The results from Isaac Gym are presented in Figure 6 and Table II, while the results from Raisim are shown in Figure 8 and Figure 9. Furthermore, to assess real-world applicability, the proposed method was evaluated in environments depicted in Figure 10, with the corresponding results summarized in Table III.
In Figure 6, the proposed method, InNKF, demonstrates a significant reduction in y-axis position error compared to InEKF. In particular, for the discrete obstacle terrain, the z-axis position error of InEKF is 0.3472m at 140s, whereas InNKF maintains an error of 0.1420m, confirming its improved performance. Table II further evaluates the performance using Absolute Trajectory Error (ATE), including NN-based state estimation as a baseline. The results indicate that InNKF consistently reduces rotation ATE across all terrains. Specifically, velocity ATE is reduced by 40.1% in the slope terrain, while position ATE is reduced by 87.0% in the discrete obstacle terrain. These results demonstrate that NN-based state estimation alone does not provide reliable performance, whereas the proposed InNKF achieves superior accuracy in most cases.
@@
InEKF NN InNKF
(MB-only) (LB-only) (Proposed)
Stairs ATER [rad] 0.1247 0.4695 0.0986
ATEv [m/s] 0.1441 0.2712 0.2056
ATEp [m] 0.6224 1.9160 0.8697
Slope ATER [rad] 0.0454 0.1212 0.0231
ATEv [m/s] 0.0930 0.0648 0.0373
ATEp [m] 1.0439 0.2363 0.1455
Uneven Terrain ATER [rad] 0.0388 0.3114 0.0341
ATEv [m/s] 0.0930 0.1744 0.1186
ATEp [m] 1.0401 1.4215 0.4582
Discrete Obstacle ATER [rad] 0.0500 0.1110 0.0215
ATEv [m/s] 0.0958 0.0631 0.0462
ATEp [m] 1.0734 0.2036 0.1389
TABLE II: Comparison of Absolute Trajectory Error (ATE) in state estimation among the two baselines and InNKF (proposed method) across different terrains.
@@
To validate its performance in a different dynamic simulator, position error was evaluated in Raisim, with additional data collected for fine-tuning. As shown in Figure 8, when the robot completed two laps along a 20m $\times$ 20m square trajectory, the x and y-axis errors between the starting and ending points remained minimal. Additionally, InNKF demonstrated significantly improved accuracy along the z-axis compared to InEKF, staying closer to the ground truth. Figure 9 compares the relative error (RE) across four environments—indoor, field, mountain, and hill—over a total walking distance of 160m. NN-based estimation exhibited high variance, reaching a maximum of 4.6298 in the indoor environment, whereas InNKF maintained a significantly lower maximum variance of 0.0466, highlighting the unreliability of NN-based estimation. As the distance traveled increased, InNKF consistently maintained a lower RE than InEKF, achieving up to 68.32% performance improvement.
@@
To verify the applicability of the proposed method in real-world environments, experiments were conducted in the setting shown in Figure 10. Since a motion capture system was unavailable, pseudo ground truth was obtained using LIO-SAM [22] with VectorNav VN-100 IMU and Ouster OS1-64-U LiDAR. Based on this, the state estimation performance was evaluated by comparing ATE and Relative Error (RE), and the results are summarized in Table III. The proposed InNKF demonstrated improvements of up to 77% in ATE, confirming its feasibility for real-world applications.
@@
InEKF NN InNKF
(MB-only) (LB-only) (Proposed)
ATER [rad] 0.2564 0.1453 0.0330
ATEp [m] 0.2042 0.1291 0.0726
RER [rad/m] 0.0210 0.0220 0.0209
(0.0021) (0.0114) (0.0021)
REp [%] 1.0019 3.3852 1.1690
(0.1727) (2.0124) (0.4459)

### v2

@@ Abstract
This paper presents an algorithm to improve state estimation for legged robots. Among existing model-based state estimation methods for legged robots, the contact-aided invariant extended Kalman filter defines the state on a Lie group to preserve invariance, thereby significantly accelerating convergence. It achieves more accurate state estimation by leveraging contact information as measurements for the update step. However, when the model exhibits strong nonlinearity, the estimation accuracy decreases. Such nonlinearities can cause initial errors to accumulate and lead to large drifts over time. To address this issue, we propose compensating for errors by augmenting the Kalman filter with an artificial neural network serving as a nonlinear function approximator. Furthermore, we design this neural network to respect the Lie group structure to ensure invariance, resulting in our proposed Invariant Neural-Augmented Kalman Filter (InNKF). The proposed algorithm offers improved state estimation performance by combining the strengths of model-based and learning-based approaches. Project webpage: https://seokju-lee.github.io/innkf_webpage
@@ III-A SE2(3) Group Generation Network
$\textbf{X}_{t}=\begin{bmatrix}\textbf{R}_{t}&v_{t}&p_{t}\\ \textbf{0}_{1,3}&1&0\\ \textbf{0}_{1,3}&0&1\\ \end{bmatrix}\in\text{SE}_{2}(3)$ (9)
@@
Using the linear combination of the generators, the neural compensator, as shown in Fig. 1, consists of three main steps. First, a temporal convolutional network (TCN) [18] takes as input a time-series sequence that includes linear acceleration, angular velocity, joint position, joint velocity, foot position, foot velocity, contact state, and the estimated state from the InEKF over the previous 49 time-steps as well as the current time step, and it produces 9 outputs. Second, these outputs are combined linearly with the generators of $\mathfrak{se}_{2}$(3) to obtain an element of $\mathfrak{se}_{2}$(3), denoted as $\hat{\eta}_{t}^{r}$. Finally, by applying the exponential mapping described in (11), the result is mapped to an element of the SE2(3) group, $\hat{\textbf{E}}_{t}^{r}$, which represents the estimated error. This value acts as a compensation term for the estimated state in the InEKF.
@@ III-B Invariant Neural-Augmented Kalman Filter
The Neural Compensator introduced in Section III-A is augmented into the InEKF [5] to form the Invariant Neural-Augmented Kalman Filter (InNKF). The structure of InNKF is shown in Fig. 2, where the state estimation is first performed using a model-based approach. In this process, two types of states are defined. The state of the InEKF is expressed as $\textbf{Z}_{t}\in\text{SE}_{N+2}(3)$, as shown in (12), while the state of the InNKF is expressed as $\textbf{X}_{t}\in\text{SE}_{2}(3)$. Here, $\textbf{X}_{t}$ is a $5\times 5$ submatrix in the top-left corner of $\textbf{Z}_{t}$.
@@
Using the SE2(3) Group Generation Network (SEGGN) designed in Section III-A, a Neural Compensator (NC) is added. The NC is trained on a dataset obtained from a simulation in which a single trajectory traverses four different terrains (stairs, slope, random uniform, discrete obstacle) for 100 seconds. The dataset included the ground truth state $\textbf{X}_{t}$, sensor data such as linear acceleration ($\tilde{a}_{t}$), angular velocity ($\tilde{\omega}_{t}$), joint position and velocity ($\tilde{q}t,\tilde{\dot{q}}t$), foot position and velocity ($\tilde{\textbf{p}}_{WC_{i}}(t),\tilde{\textbf{v}}_{WC_{i}}(t)$), contact state ($\text{S}_{t}$), and the estimated state $\bar{\textbf{X}}_{t}^{+}$ from the InEKF. This dataset was used to train the SEGGN through the supervised learning process shown in Fig. 3.
@@ IV-A Performance Analysis of SE2(3) Group Generation
The neural network used in this study does not perform operations directly on the Lie group but operates in the Lie algebra, followed by exponential mapping. Since the Lie group is defined on a smooth manifold, performing group operations requires matrix multiplication. In contrast, the Lie algebra, defined in the tangent space, only requires matrix addition for operations. Consequently, obtaining the output in the Lie algebra and applying exponential mapping significantly improves training efficiency. Table I and Fig. 4 show the computation time according to batch size. While matrix addition has a time complexity of O$(n^{2})$, matrix multiplication has a time complexity of O$(n^{3})$. This highlights the advantage of obtaining outputs in the tangent space, where only matrix addition is required, to enhance training efficiency. A difference of approximately 0.1 times in computation time was observed, confirming significantly faster performance.
Furthermore, due to the structure of the SEGGN, which utilizes the bases of $\mathfrak{se}_{2}(3)$ followed by an exponential mapping process, it inevitably produces outputs within SE2(3). The corresponding results are presented in Fig. 5. The first row of Fig. 5 illustrates the transformation of the original coordinates based on the output values obtained from SEGGN. Visually, it can be observed that the transformed coordinates maintain orthonormality, ensuring the structural integrity of the transformation. A numerical analysis of these results is presented in the histograms below. The Frobenius norm error distribution shows that nearly all samples have errors close to zero, indicating minimal deviation from the expected transformation properties. Additionally, the determinant values are consistently close to one, confirming that the rotational component of the output matrix adheres well to the properties of the SO(3) group (26).
@@ IV-B Implementation Details
Figure 6: Position estimation results across four terrains (stairs, slope, uneven terrain, discrete obstacle) over a time interval of 0 $\sim$ 160 s. The solid black line represents the ground truth, the blue dashed line represents InEKF, and the red dashed line represents the InNKF (proposed method).
To reduce the sim-to-real gap, sensor noise and IMU bias were introduced to reflect real-world conditions better. These noise factors were applied during both locomotion controller training and dataset collection. The sensor noise was assumed to follow a Gaussian distribution. Additionally, a covariance matrix was configured to address uncertainties in the measurement model. Since InEKF measurements primarily depend on kinematic information, the covariance matrix was constructed based on the standard deviation of the noise applied to joint position and velocity.
@@ IV-C Experimental Results
Three baselines were compared to analyze the experimental results. The first was the model-based (MB) estimator, represented by the Invariant Extended Kalman Filter (InEKF). The second was the learning-based (LB) estimator, represented by a Neural Network (NN), and the third was the hybrid estimator, represented by KalmanNet.
The training and test datasets were collected in the Isaac Gym simulator [19], using terrains shown in Fig. 7. The quadruped robot, Unitree Go1, was the model used, controlled by a blind locomotion controller trained via reinforcement learning [20]. To evaluate whether the state estimator overfits the training conditions, additional experiments were conducted in Raisim [21]. The results from Isaac Gym are presented in Fig. 6 and Table II, while the results from Raisim are shown in Fig. 8 and Fig. 9. Furthermore, to assess real-world applicability, the proposed method was evaluated in environments depicted in Fig. 10, with the corresponding results summarized in Table III.
In Fig. 6, the proposed method, InNKF, demonstrates a significant reduction in y-axis position error compared to InEKF. In particular, for the discrete obstacle terrain, the z-axis position error of InEKF is 0.3472m at 160s, whereas InNKF maintains an error of 0.1420m, confirming its improved performance. Table II further evaluates the performance using Absolute Trajectory Error (ATE), including NN-based state estimation and KalmanNet which is the hybrid state estimation as the baselines. The results indicate that InNKF consistently reduces rotation ATE across all terrains. Specifically, velocity ATE is reduced by 40.1% in the slope terrain, while position ATE is reduced by 87.0% in the discrete obstacle terrain. These results demonstrate that NN-based state estimation alone does not provide reliable performance, whereas the proposed InNKF achieves superior accuracy in most cases. Also, the hybrid method KalmanNet showed good performance in rotation and position on the stairs terrain, but in most cases, it did not outperform InNKF.
@@
InEKF NN KalmanNet InNKF
(MB-only) (LB-only) (Hybrid) (Proposed)
Stairs ATER [rad] 0.1247 0.4695 0.0809 0.0986
ATEv [m/s] 0.1441 0.2712 0.2579 0.2056
ATEp [m] 0.6224 1.9160 0.3991 0.8697
Slope ATER [rad] 0.0454 0.1212 0.0591 0.0231
ATEv [m/s] 0.0930 0.0648 0.0717 0.0373
ATEp [m] 1.0439 0.2363 0.5644 0.1455
Uneven Terrain ATER [rad] 0.0388 0.3114 0.0502 0.0341
ATEv [m/s] 0.0930 0.1744 0.3004 0.1186
ATEp [m] 1.0401 1.4215 0.5834 0.4582
Discrete Obstacle ATER [rad] 0.0500 0.1110 0.0732 0.0215
ATEv [m/s] 0.0958 0.0631 0.0263 0.0462
ATEp [m] 1.0734 0.2036 0.6053 0.1389
TABLE II: Comparison of Absolute Trajectory Error (ATE) in state estimation among the three baselines and InNKF (proposed method) across different terrains.
@@
To validate its performance in a different dynamic simulator, position error was evaluated in Raisim, with additional data collected for fine-tuning. As shown in Fig. 8, when the robot completed two laps along a 20m $\times$ 20m square trajectory, the x and y-axis errors between the starting and ending points remained minimal. Additionally, InNKF demonstrated significantly improved accuracy along the z-axis compared to InEKF, staying closer to the ground truth. Fig. 9 compares the relative error (RE) across four environments—indoor, field, mountain, and hill—over a total walking distance of 160m. NN-based estimation exhibited high variance, reaching a maximum of 4.6298 in the indoor environment, whereas InNKF maintained a significantly lower maximum variance of 0.0466, highlighting the unreliability of NN-based estimation. As the distance traveled increased, InNKF consistently maintained a lower RE than InEKF, achieving up to 68.32% performance improvement.
@@
To verify the applicability of the proposed method in real-world environments, experiments were conducted in the setting shown in Fig. 10. Since a motion capture system was unavailable, pseudo ground truth was obtained using LIO-SAM [22] with VectorNav VN-100 IMU and Ouster OS1-64-U LiDAR. Based on this, the state estimation performance was evaluated by comparing ATE and Relative Error (RE), and the results are summarized in Table III. The proposed InNKF demonstrated improvements of up to 77% in ATE, confirming its feasibility for real-world applications.
@@
InEKF NN KalmanNet InNKF
(MB-only) (LB-only) (Hybrid) (Proposed)
ATER [rad] 0.2564 0.1453 0.1627 0.0330
ATEp [m] 0.2042 0.1291 0.2166 0.0726
RER [rad/m] 0.0210 0.0220 0.0211 0.0209
(0.0021) (0.0114) (0.0043) (0.0021)
REp [%] 1.0019 3.3852 1.2109 1.1690
(0.1727) (2.0124) (0.2729) (0.4459)